# Estimating Player Performance in Different Contexts Using Fine-tuned Large Events Models

### v1

@@ 4. Experiments
4.1. Simulating the Premier League
@@
Table 4. Comparing the forecasts from the finetuned LEM models against the end-of-season (EoS) tables. The average displacement (Avg. Disp.) is the average number of team positions from the actual position. We compare against the full and home tables, with the home table providing less displacement since all models are finetuned using home game data.
@@
An indepth analysis of the stats generated by fine-tuned models is presented in Annex A.
4.2. Cristiano Ronaldo vs Lionel Messi
In Figures 5 and 6, we explore the hypothetical influence of football superstars Cristiano Ronaldo and Lionel Messi on the home performance of teams. Utilizing fine-tuned predictive models, the violin plots reveal the distribution of expected home points with the addition of either player to the teams. These distributions allow us to compare the mean, variability, and shape of the teams’ points distribution under the baseline conditions and the two alternative scenarios.
Figure 5. The expected impact of adding Cristiano Ronaldo or Lionel Messi on the teams in the EPL. The figure presents the violin plots of the simulations using the fine-tuned models.
Figure 6. The expected impact of replacing a player for Cristiano Ronaldo or Lionel Messi on the teams in the EPL.
The ”Team” data, indicated by the black and white violin, shows the distribution of points without the influence of star players. There is a noticeable variation between teams, with some, like Manchester City and Liverpool, showing a higher median and denser distribution towards the top end of the scale, suggesting consistently high performance at home games.
@@ 4.3. The importance of context
Figure 7. The impact on the baseline of the top 10 players of the 2017/2018 season, according to Sofascore.
Figure 8. The expected impact of adding Casemiro or Illarramendi on the teams in the EPL.
Figures 7 and 8 have the role of explaining that context is everything. Figure 7 presents the expected impact of players if they were placed in the original context of the LEM, i.e., placed in a team that is the average of all teams. Figure 9 shows why it is wrong to use this general context. We observe that, when introducing the best (Illarramendi) and the worst (Casemiro) performers from Figure 7 in the teams in the Premier League, the difference between them is smaller than first suggested. While Illarramendi still outperforms Casemiro most times, the difference between the impact on the teams is lower than expected.
5. Discussion
@@ Appendix A Stats from Simulations
Table 5. Per game average of several team statistics in our simulations. The ’Home’ columns detail the actions performed by the team, while the ’Away’ columns represent the actions performed by the opposition teams

### v2

@@ 4 Experiments
4.1 Simulating the Premier League
@@
Table 5: Comparing the forecasts from the finetuned LEM models against the end-of-season (EoS) tables. The average displacement (Avg. Disp.) is the average number of team positions from the actual position. We compare against the full and home tables, with the home table providing less displacement since all models are finetuned using home game data.
@@
An in-depth analysis of the stats generated by fine-tuned models is presented in Annex A.
4.2 Cristiano Ronaldo vs Lionel Messi
In Figures 4 and 5, we explore the hypothetical influence of football superstars Cristiano Ronaldo and Lionel Messi on the home performance of teams. Utilizing fine-tuned predictive models, the violin plots reveal the distribution of expected home points with the addition of either player to the teams. These distributions allow us to compare the mean, variability, and shape of the teams’ points distribution under the baseline conditions and the two alternative scenarios.
Figure 4: The expected impact of adding Cristiano Ronaldo or Lionel Messi on the teams in the EPL. The figure presents the violin plots of the simulations using the fine-tuned models.
Figure 5: The expected impact of replacing a player for Cristiano Ronaldo or Lionel Messi on the teams in the EPL.
The "Team" data, indicated by the black and white violin, shows the distribution of points without the influence of star players. There is a noticeable variation between teams, with some, like Manchester City and Liverpool, showing a higher median and denser distribution towards the top end of the scale, suggesting consistently high performance at home games.
@@ 4.3 The importance of context
Figure 6: The impact on the baseline of the top 10 players of the 2017/2018 season, according to Sofascore.
Figure 7: The expected impact of adding Casemiro or Illarramendi on the teams in the EPL.
Figures 6 and 7 have the role of explaining that context is everything. Figure 6 presents the expected impact of players if they were placed in the original context of the LEM, i.e., placed in a team that is the average of all teams. Figure 9 shows why it is wrong to use this general context. We observe that, when introducing the best (Illarramendi) and the worst (Casemiro) performers from Figure 6 in the teams in the Premier League, the difference between them is smaller than first suggested. While Illarramendi still outperforms Casemiro most times, the difference between the impact on the teams is lower than expected.
5 Discussion
@@ Appendix A Stats from Simulations
Table 6: Per game average of several team statistics in our simulations. The ’Home’ columns detail the actions performed by the team, while the ’Away’ columns represent the actions performed by the opposition teams